OpenStudy (anonymous):

if y varies inversely as x and y=5 when x=2 find x when y is 4. x=? help

OpenStudy (anonymous):

because y varies inversely as x,then y=k/x,where k is a constant. when x=2,5=k/2,k=10 y=10/x when y=4, 4=10/x 4x=10 x=10/4=5/2